# Cole: A Hierarchical Generation Framework for Graphic Design

### 28 Nov 2023 v1

@@ Abstract
Graphic design, which has been evolving since the 15th century, plays a crucial role in advertising. The creation of high-quality designs demands creativity, innovation, and lateral thinking. This intricate task involves understanding the objective, crafting visual elements such as the background, decoration, font, color, and shape, formulating diverse professional layouts, and adhering to fundamental visual design principles. In this paper, we introduce Cole 111In 1843, Henry Cole introduced the world’s first commercial Christmas card [37]. We commemorate his contributions with the Cole system., a hierarchical generation framework purposed to comprehensively address these challenges. This Cole system can transform a straightforward intention prompt into a high-quality graphic design, while also supporting flexible editing based on user input. Examples of such input might include directives like “design a poster for Hisaishi’s concert.” The key insight is to dissect the complex task of text-to-design generation into a hierarchy of simpler sub-tasks, each addressed by specialized models working collaboratively. The results from these models are then consolidated to produce a cohesive final output. Our hierarchical task decomposition can streamline the complex process and significantly enhance generation reliability. Our Cole system consists of multiple fine-tuned Large Language Models (LLMs), Large Multimodal Models (LMMs), and Diffusion Models (DMs), each specifically tailored for a design-aware text or image generation task. Furthermore, we construct the DesignerIntention benchmark to highlight the superiority of our Cole over existing methods in generating high-quality graphic designs from user intent. We perceive our Cole as an important step towards addressing more complex visual design generation tasks in the future.
@@ 1 Introduction
The recent advancement in the quality of natural image generation has been remarkable, elevating it to the level of professional photography. This leap forward can be attributed to the development of technologies such as Imagen [7], SDXL [26], and DALLE3 [2]. The pivotal factors contributing to these advancements include the use of powerful Large Language Model (LLM) as a text encoder, the scaling up of training datasets, increased model complexity, the design of advanced sampling strategies, and the enhancement of data quality, among others. We believe that now is the opportune moment to redirect our efforts towards more professional image generation, particularly in the area of graphic design generation considering its pivotal roles in advertising, branding, and marketing.
@@
The most recent study [18] has empirically shown that the groundbreaking DALLE3 exhibits impressive capabilities in generating high-quality design images, marked by visually striking graphics and layouts, as illustrated in Figure 2. However, these images are not without their flaws. They continue to confront several critical challenges, such as mis-rendered visual text, often resulting in missing or extra characters, a phenomenon also highlighted in [2]. Furthermore, editing these generated images demands complex operations such as segmentation, erasing, inpainting, and others, owing to their inherently uneditable nature. Another substantial limitation is the necessity for users to provide detailed text prompts - a task in graphic design generation that typically requires a high degree of professional expertise to develop effective prompts. In contrast to DALLE3, our Cole system, as demonstrated in Figure 1, is capable of generating superior quality graphic design images while only necessitating simple user intention.
@@
We maintain that these three limitations critically undermine the quality of the generated graphic design images. A scalable, high-quality graphic design generation system should ideally require minimal effort from users, produce accurate and high-quality typography information for a variety of purposes, and offer a flexible editing space. This would allow users to further refine the output, integrating human expertise when necessary. Our objective in this work is to establish an efficient and dependable autonomous text-to-design system capable of generating high-quality graphic design images from user intent prompts. We suggest breaking down the complex process of graphic design image generation into a hierarchical generation process. This process involves multiple specialized generation models, each designed to handle different sub-tasks:
First, the focus is on creative design and interpretation, essentially focusing on understanding intentions. This is accomplished by employing state-of-the-art LLMs, specifically the Llama-B [33] and fine-tuning it with a comprehensive dataset of nearly curated intention-JSON pairs. The JSON file contains critical design-related data, such as background captions, object captions, and textual details. Additionally, we provide optional specifications for object positioning and others.
Second, our focus shifts to visual composition and enhancement, encompassing two sub-tasks: the generation of visual elements and the creation of typography attributes. This involves fine-tuning specialized cascaded diffusion models like DeepFloyd/IF [19] to generate diverse visual elements. The construction of these models ensures the harmonious integration of elements like the embellished background and object image layers. We then use a typography Large Multimodal Model (LMM), built with LLaVA--B [24], to predict the typography JSON file. This leverages the Design LLM’s forecasted JSON file, a diffusion model’s projected background image, and a cascaded diffusion model’s anticipated object image. A graphic renderer then assembles these elements according to the layout in the predicted JSON file.
Third, the process concludes with feedback and quality assurance, aimed at enhancing the overall design quality. This involves meticulously fine-tuning a reflection LMM and employing GPT-4V(ision) for a thorough, multi-faceted quality analysis. This final step facilitates the necessary refinement of the JSON file, including adjustments to the text box position and sizes as deemed necessary.
@@
Last, to access the capability of our system, we construct a DesignerIntention, which consists of professional graphic design intention prompts covering diverse categories and creative ones, to compare our method with existing state-of-the-art image generation system, conduct thorough ablation experiments for each generation model on different sub-tasks, offer a detailed analysis of the graphic designs created by our system, and engage in a discussion about both the limitations and prospective future pathways of graphic design image generation.
@@ 2 Related Work
Autonomous Agent based on LLMs Recent studies have proposed the development of advanced agent systems leveraging powerful Large Language Models (LLMs) to tackle a variety of challenging tasks. We can classify these into two primary categories: (i) General Multi-Agent Systems: These attempts aim to create an agent capable of addressing as many tasks as possible, ranging from coding, mathematics, science to societal related tasks. Among these attempts, several studies [11, 27, 40, 6, 40, 22] have explored the similar communicative (or cooperative) multi-agent framework to deconstruct a complex task into multiple sub-tasks, each handled by a specialized agent. For instance, ChatDev [27] introduces a chat-based software development framework to break down the development process into sequential atomic subtasks, where each subtask demands collaborative interaction and cross-examination between two roles. (ii) Image Generation Agents: These studies [45, 20, 39, 44, 29] concentrate more on image generation. For example, the concurrent work Idea2Img [45] unveils a framework for automated image generation, supporting complex prompts, and enhancing any text-to-image model for image design. Mini DALLE3 [20], a DALLE3 derivative, provides an interactive text-to-image generation system for a seamless user experience.
Our model primarily distinguishes itself from existing image generation agent studies [45, 20] and multi-agent efforts [27, 40, 22] in the mechanism of task decomposition and the usage of tools. While most existing studies incorporate tool usage knowledge by designing tailored prompt instructions or in-context examples for GPT-/, or GPT-4V(ision) to iteratively refine image generation prompts, we empirically demonstrate that these approaches fall short in excelling at professional graphic design image generation tasks due to the complexity of the required expertise. Unlike these efforts, our Cole decomposes the complicated graphic design image generation task into a hierarchical set of generation tasks, and performs dedicated fine-tuning of each generation model on our constructed professional structured graphic design dataset.
Design Image Generation Graphic design generation, given its extensive applications in advertising, branding, and marketing, is a vital research area with substantial industry impact. As discussed in [13, 9, 12, 21, 8, 23, 4, 46], a well-crafted graphic design can effectively convey messages through a skillful blend of shapes, colors, and fonts, all while respecting design principles and aesthetics. A significant amount of research has been invested in generating specific components of professional graphic design images, such as layouts or typography attributes. For instance, FlexDM [16, 43] leverages a unified Transformer architecture for element prediction. Additional research [30] focuses on typography generation, while UI-Diffuser [35] investigates mobile UI generation. On the other hand, the majority of studies [12, 23, 4, 46] primarily concentrate on layout generation, assuming that all other information is at hand.
Unlike the above-mentioned methods that focus on individual aspects within the graphic design generation task, this study aims to develop an autonomous, comprehensive system capable of executing all design generation tasks from clear user intent descriptions. This task formulation presents a significantly greater challenge compared to most existing approaches. To the best of our knowledge, our Cole is among the pioneering works that seek to tackle the complex task of autonomous graphic design generation based solely on a basic user intent prompt.
@@ 3.1 Cole Framework Overview
Motivated by the necessity to generate high-quality graphic designs that accurately reflect user intentions, we recognize the importance of a seamless integration of multiple roles. Each of these roles is designed to address specific aspects of the graphic design process, as indicated in the references [36, 34]. Consequently, we have broken down the text-to-design generation task into a hierarchical structure of tasks. These consist of the intention-to-JSON generation task, text-to-background image generation task, text-to-object image and alpha mask generation task, typography attribute generation task, and the graphic rendering task. The tasks of intention-to-JSON generation and typography attribute generation both follow a hierarchical structure. They necessitate the prediction of missing values given a predefined structured request.
@@
We introduce the comprehensive hierarchical framework of our Cole system as depicted in Figure 3. To effectively train the specialized design-agent system, we initially amass approximately 100,000 high-quality raw graphic design images from the internet. Following this, we organize these graphic design data into a hierarchical structure which includes components such as (intention, JSON) pairs, (caption, background-image) pairs, (caption, object-image) pairs, and (text-image, typography JSON) pairs. We will further explore the process of constructing each component in the ensuing discussion.
3.2 Design LLM
We first formulate the intention-to-JSON prediction task as a masked field prediction problem following [15]. This approach learns to predict element types, positions, attributes, and text using a unified Transformer architecture. Furthermore, we adopt a strategy from [1] to construct causally masked multi-modal LLMs by leveraging the Llama-B [33].
@@ 3.2 Design LLM
We have fine-tuned the causally masked Llama-B model using our assembled dataset of nearly (intention, JSON) pairs, over an approximate span of epochs. Our empirical findings suggest that incorporating additional predictions of the complete caption, along with key words from the graphic design image, facilitates subsequent tasks of background and object image generation.
Given that both intentions and detailed captions are absent in the raw graphic design data, we first employ GPT- and GPT- models to generate the intentions. This is accomplished with the aid of the system prompt displayed in the “Intention Generation Prompt”. For a more comprehensive understanding, we have included detailed intention generation prompt with in-context examples in the supplementary materials. Subsequently, we utilize the recently developed, large multimodal model known as LLaVA--B model [24]. This model is effective in generating detailed captions for the entire image, background image, and object image captions.
Figure 4 presents a representative example, showcasing the ground-truth triplets of graphic design images, alongside the generated user intentions and corresponding JSON files. In addition, we have also investigated the potential of the highly proficient GPT- model for directly generating JSON from user intentions. This is achieved by formulating an instructive prompt. Our empirical analysis, detailed in the experiment section, demonstrates that our methodology consistently surpasses the performance of GPT-.
@@
3.3 Text-to-Background Diffusion Model
@@
Given the current unavailability of mature text-to-SVG technology and the scarcity of high-quality text-SVG pairs, which leads to the absence of strong text-to-SVG generation models [17, 47, 41, 42], our study suggests consolidating all SVG elements and additional embellishments into one unified image layer. Therefore, we first extract the generated detailed description of the background image layer from the JSON file. Then we train a text-to-background diffusion model to generate the coherent background images consisting of all the SVG files and decoration elements.
@@
We opt to fine-tune our collected less than (caption, background image) pairs using the recent DeepFloyd IF [19] diffusion models for epochs. The output resolution is . Following this, we employ the generated background images as conditioning input for the text-to-object diffusion model, which predicts coherent high-resolution object image layers based on their corresponding object prompts.
@@
3.4 Text-to-Object Diffusion Model

### 18 Mar 2024 v2

# Cole: A Hierarchical Generation Framework for Multi-Layered and Editable Graphic Design

@@ Abstract
Graphic design, which has been evolving since the 15th century, plays a crucial role in advertising. The creation of high-quality designs demands design-oriented planning, reasoning, and layer-wise generation. This intricate task involves understanding vague intentions and faithfully generating multi-layered visual elements, including the background, decorations, objects, and fonts, among others. It also requires layout planning for all elements and visual reasoning to generate all visual elements that satisfy visual design principles. Unlike the recent CanvaGPT, which integrates GPT-4 with existing design templates to build a custom GPT, this paper introduces the Cole 111In 1843, Henry Cole introduced the world’s first commercial Christmas card [36]. We commemorate his contributions with the Cole system. system—a hierarchical generation framework designed to comprehensively address these challenges. This Cole system can transform a vague intention prompt into a high-quality multi-layered graphic design, while also supporting flexible editing based on user input. Examples of such input might include directives like “design a poster for Hisaishi’s concert.” The key insight is to dissect the complex task of text-to-design generation into a hierarchy of simpler sub-tasks, each addressed by specialized models working collaboratively. The results from these models are then consolidated to produce a cohesive final output. Our hierarchical task decomposition can streamline the complex process and significantly enhance generation reliability. Our Cole system comprises multiple fine-tuned Large Language Models (LLMs), Large Multimodal Models (LMMs), and Diffusion Models (DMs), each specifically tailored for design-aware layer-wise captioning, layout planning, reasoning, and the task of generating images and text. Furthermore, we construct the DesignerIntention benchmark to demonstrate the superiority of our Cole system over existing methods in generating high-quality graphic designs from user intent. Last, we present a Canva-like multi-layered image editing tool to support flexible editing of the generated multi-layered graphic design images. We perceive our Cole system as an important step towards addressing more complex and multi-layered graphic design generation tasks in the future.
@@ 1 Introduction
The recent advancement in the quality of natural image generation has been remarkable, elevating it to the level of professional photography. This leap forward can be attributed to the development of technologies such as Imagen [8], SDXL [26], and DALLE3 [2]. The pivotal factors contributing to these advancements include the use of powerful Large Language Model (LLM) as a text encoder, the scaling up of training datasets, increased model complexity, the design of advanced sampling strategies, and the enhancement of data quality, among others. We believe that now is the opportune moment to redirect our efforts towards more professional image generation, particularly the graphic design generation considering its pivotal roles in advertising, branding, and marketing.
@@
The most recent study [17] has empirically shown that the groundbreaking DALLE3 exhibits impressive capabilities in generating high-quality design images, marked by visually striking graphics and layouts, as illustrated in Figure 2. However, these images are not without their flaws. They continue to confront several critical challenges, such as mis-rendered visual text, often resulting in missing or extra characters, a phenomenon also highlighted in [2]. Furthermore, editing these generated images demands complex operations such as segmentation, erasing, inpainting, and others, owing to their inherently uneditable nature. Another substantial limitation is the necessity for users to provide detailed text prompts - a task in graphic design generation that typically requires a high degree of professional expertise to develop effective prompts. In contrast to DALLE3, our Cole system, as demonstrated in Figure 1, is capable of generating multi-layered and editable graphic design images from simple user intentions.
The most recent study [17] has empirically shown that the groundbreaking DALLE3 exhibits impressive capabilities in generating high-quality design images, marked by visually striking graphics and layouts, as illustrated in Figure 2. However, these images are not without their flaws. They continue to confront several critical challenges, such as mis-rendered visual text, often resulting in missing or extra characters, a phenomenon also highlighted in [2]. Furthermore, editing these generated images demands complex operations such as segmentation, erasing, inpainting, and others, owing to their inherently uneditable nature. Another substantial limitation is the necessity for users to provide detailed text prompts - a task in graphic design generation that typically requires a high degree of professional expertise to develop effective prompts. In contrast to DALLE3, our Cole system, as demonstrated in Figure 1, is capable of generating multi-layered and editable graphic design images from simple user intentions.
@@
We maintain that these three limitations critically undermine the quality of the generated graphic design images. A scalable, high-quality graphic design generation system should ideally require minimal effort from users, produce accurate and high-quality typography information for a variety of purposes, and offer a flexible editing space. This would allow users to further refine the output, integrating human expertise when necessary.
| Method | Accuracy visual text | Multi-layered output | Flexible editing | Generated design | GPT4-V Score | ||||
| Layout | Image | Typography | Content | Innovation | |||||
| DeepFloyd/IF [18] | ✗ | ✗ | ✗ | ✓ | 65.31 | 56.59 | 55.32 | 83.69 | 78.23 |
| SDXL [26] | ✗ | ✗ | ✗ | ✓ | 60.89 | 60.12 | 61.21 | 67.33 | 69.92 |
| DALLE3 [2] | ✗ | ✗ | ✗ | ✓ | 80.29 | 69.77 | 74.73 | 89.16 | 90.61 |
| CanvaGPT [4] | ✓ | ✓ | ✓ | ✗ | 75.66 | 57.01 | 75.35 | 70.10 | 63.27 |
| Ours Cole | ✓ | ✓ | ✓ | ✓ | 80.82 | 72.53 | 75.46 | 87.44 | 88.72 |
Our objective in this work is to establish an efficient and dependable autonomous text-to-design system capable of generating high-quality graphic design images from user intent prompts. We suggest breaking down the complex process of graphic design image generation into a hierarchical generation process. This process involves multiple specialized generation models, each designed to handle different sub-tasks: First, we train a Design-LLM to comprehend vague intentions and undertake design-oriented planning tasks. We formulate this as a text-to-JSON prediction task, which requires transforming naive and sparse user intentions into structured, layer-wise captions or attributes essential for the following task decomposition. Second, to perform design-oriented visual reasoning and generation, we train additional three models including (i) a text-to-background diffusion model for background layer generation: it is crucial to guarantee that the generated background images reserve adequate space for the incorporation of various elements, which is different from conventional methods such as DALLE3, which often produce objects filling the entire image space. (ii) a text-to-object diffusion model for object layer generation, it’s essential to ensure the model to understand both the content and style of the background image. Subsequently, the model should generate an object layer that not only aligns stylistically but is also placed in logical positions. (iii) a Typography-LMM for typography prediction, it needs to predict distinct typography attributes for each text box. This component requires a sophisticated level of reasoning about the visual contents of the merged background and object image layers, then predicting the typography attributes accordingly. We also demonstrate the generalization capability of our Typography-LLM by applying it to processed design images originally generated using DALLE3. Refer to Figure 1 for the visual results. Besides, we also introduce a feedback and quality assurance process to enhance the quality.
@@
Last, to access the capability of our system, we construct a DesignerIntention, which consists of professional graphic design intention prompts covering diverse categories and creative ones, to compare our method with existing state-of-the-art image generation system, conduct thorough ablation experiments for each generation model on different sub-tasks, offer a detailed analysis of the graphic designs created by our system, and engage in a discussion about both the limitations and prospective future pathways of graphic design image generation. Figure 3 shows that, among both non-designers and designers, our Cole system outperforms DALLE3 in text fidelity and message conveyance, achieving win rates of 69.6% and 70.5%, respectively. Regarding visual appeal, non-designers show a slight preference for Cole, while designers favor DALLE3. Our results are also preferred by designers over those from the latest commercial product, CanvaGPT. Table 1 provides a detailed comparison highlighting the key differences between our approach, earlier text-to-image generation models, and the latest CanvaGPT. Additionally, we present GPT4-V scores from various perspectives, with our Cole outperforming others in three critical areas: layout, image, and typography quality.
@@ 2 Related Work
Autonomous Agent based on LLMs Recent studies [44, 19, 38, 43, 28, 11, 27, 39, 7, 39, 21] have proposed the development of advanced (multi-)agent systems leveraging powerful LLMs to tackle various tasks. Some of them [44, 19, 38, 43, 28] concentrate more on image generation tasks. Our model primarily distinguishes itself from these works in the mechanism of task decomposition and the usage of tools. While most existing studies incorporate tool usage knowledge by designing tailored prompt instructions or in-context examples for GPT-/, or GPT-4V(ision) to iteratively refine the text prompts, our Cole decomposes the complicated graphic design image generation task into a hierarchical set of generation tasks, and performs dedicated fine-tuning of each generation model on our constructed professional graphic design dataset.
Design Image Generation The majority of studies [12, 22, 5, 45, 29, 34, 15, 42] primarily concentrate on individual aspects within the graphic design generation task, assuming that all other information is at hand. Unlike the these methods, this study aims to develop an autonomous, comprehensive system capable of executing all design generation tasks from clear user intent descriptions. This task formulation presents a significantly greater challenge compared to most existing approaches. To the best of our knowledge, our Cole is among the pioneering works that seek to tackle the complex task of autonomous graphic design generation based solely on a basic user intent prompt.
@@ 3.1 Cole Framework Overview
Motivated by the necessity to generate high-quality graphic designs that accurately reflect user intentions, we recognize the importance of a seamless integration of multiple roles. Each of these roles is designed to address specific aspects of the graphic design process, as indicated in the references [35, 33]. Consequently, we train an LLM for multi-layered task planning, breaking down the text-to-design generation into a hierarchical structure of tasks. These include the intention-to-JSON generation, text-to-background image generation, text-to-object image and alpha mask generation, typography attribute generation, and the multi-layered graphic editing and rendering tasks. The tasks of intention-to-JSON and typography attribute generation both adhere to a hierarchical structure, requiring the prediction of missing values based on a predefined structured request.
@@
We introduce the comprehensive hierarchical framework of our Cole system, as depicted in Figure 4. To effectively train the Cole system, we initially amass approximately 100,000 high-quality raw graphic design images from the internet, covering multiple graphic design sources. Following this, we organize the graphic design data into a hierarchical structure that includes components such as (intention, JSON) pairs, (caption, background-image) pairs, (caption, object-image) pairs, and (text-image, typography JSON) pairs. We will further explore the process of constructing each component in the ensuing discussion.
3.2 Design LLM: Intention Recaption and Layout Planning
We first detail the process of generating (intention, JSON) pair data for graphic design and then demonstrate the implementation of the Design LLM, aimed at intention recapitulation and planning multi-layered layouts.
Creating Intention for Design Image Given that both intentions and detailed captions are absent in the raw graphic design data, we first employ GPT- to generate the intentions from the raw image information. This includes the image title, format, keywords, and all visual text present in the image. This is accomplished with the aid of the system prompt displayed in the “Intention Generation Prompt” (shown in Supplementary). The left of Figure 5 presents a representative example, showcasing the ground-truth pairs of graphic design images and the generated user intentions. We generate the intention for nearly graphic design images.
Creating JSON for Design Image Subsequently, we prepare the structured JSON data to represent the design images, which include multiple key-value pairs such as ‘global caption’, ‘category’, ‘keywords’, and ‘layers’. Within ‘layers’, we further decompose it into three sub-layers: ‘background layer’, ‘object layer’, and ‘text layer’. Considering that not all design images contain the object layer, we also use a ‘flag’ to mark whether the object layer exist. For all captions, we opt to utilize the recently developed large multimodal model known as the LLaVA-1.5-13B model [24], which is effective in generating detailed captions for the entire image, including background and object image captions. The right side of Figure 5 presents a representative example, showcasing the generated JSON file for the given image. We create matching JSON files for nearly graphic design images, each paired with the above generated intentions.
@@ 3.2 Design LLM: Intention Recaption and Layout Planning
Implementation and Training Settings After constructing the intention and JSON pairs for the graphic design images, we formulate the intention-to-JSON prediction task as a masked field prediction problem following [14], which is very different from the conventional text-to-text prediction task. This approach learns to translate intentions into detailed JSON files and conducts multi-layered layout planning by predicting the global captions, background layer captions, object layer captions, contents of the text layers including the heading, sub-heading, and body-text. Furthermore, we adopt a strategy from [1] to construct causally masked multi-modal LLMs by leveraging the Llama-B [32]. We have fine-tuned the causally masked Llama-B model using our assembled dataset of nearly (intention, JSON) pairs, over an approximate span of epochs. Our empirical findings suggest that incorporating additional predictions of the ‘global caption’, along with ‘keywords’ from the graphic design image, facilitates subsequent tasks of background and object image generation.
@@
3.3 Text-to-Background Diffusion Model: Visual Planning to Generate Canvas Placeholders
We now present the methodology for preparing (caption, background-image) pairs and delve into the complexities of training a dependable background image to serve as a canvas placeholder. This involves strategically planning about where to allocate sufficient space for the seamless integration of diverse elements.
@@ 3.3 Text-to-Background Diffusion Model: Visual Planning to Generate Canvas Placeholders
Preparing Text-Background Pairs Given the current unavailability of mature text-to-SVG technology and the scarcity of high-quality text-SVG pairs, which leads to the absence of strong text-to-SVG generation models [16, 46, 40, 41], our study suggests consolidating all SVG elements and additional embellishments into one unified background image layer. Then we extract the generated detailed description of the unified background image layer from the JSON file. The left side of Figure 6 displays an example of the decomposed background image layer from a given graphic design image. We construct less than (caption, background image) pairs for use as training data.
@@
Implementation and Training Settings Then we train a text-to-background diffusion model to generate the coherent background images consisting of all the SVG files and decoration elements. We opt to fine-tune our collected (caption, background image) pairs using the DeepFloyd IF [18] diffusion models for epochs. The output resolution is . Following this, we employ the generated background images as conditioning input for the text-to-object diffusion model, which predicts coherent high-resolution object image layers based on their corresponding object prompts.
One key challenge in generating the background image lies in ensuring it serves effectively as a canvas placeholder. This requires reserving adequate space for incorporating various elements, including object image layers and text layers. Consequently, the text-to-background generation model must perform visual planning to determine where to allocate space. This approach significantly diverges from conventional methods like DALLE3, which often generate objects occupying the entire image space. We provide a qualitative comparison of the background images generated by different methods in the appendix. Another observed challenge is the length of the generated detailed captions, which can extend to tens of words, resulting in hundreds of text tokens - far exceeding the original default maximum length. To overcome this, we have increased the maximum number of text tokens from to .
Offset-noise is important for background image generation We observe the graphic design background images tend to have a uniform color distribution and this aspect sets them apart from the pre-trained distributions of photorealistic images. Our empirical findings emphasize the effectiveness of the offset noise technique [23, 9] for high-quality background image generation. This is because, even when , the model’s input isn’t purely noise. The leaked signal carries low-frequency information, such as the overall mean of each channel. The implementation details are as follows: . We choose by default.
@@
3.4 Text-to-Object Diffusion Model: Visual Reasoning based on the Generated Background Image
This section outlines the method for creating (caption, object image) pair data and showcases the deployment of the text-to-object diffusion model. This model is designed to produce an object layer that is not only stylistically coherent but also logically positioned. A key technical insight in this endeavor is the model’s need to predict an additional composite image, a step that is essential for ensuring the precision of the object placements.